Use the alternating series test to prove that P (-1) 2ºn2 - converges. n!
Answer #1

This is the proper solution of your question.

Thank you.

S (-1) nn? na na Alternate series test for all n (:) but <bn (ii) limebn=0 no n-00 then cho = 2nn ni bau = 2n+ (n+1)? Dr Anti

but = 2 (nte) { when > 2 bn n² n-00 noon! og (-2)^ (2) n n=3 n! Converget os (1) (2) n? Converges -